OpenStudy (anonymous):

ok I need a real math wiz Scenario: Suppose your business makes high-grade PDAs, personal digital assistants. It costs you $80,000 a month for fixed costs, such as rent, employee salaries, and other essential items. Also, it costs $300 to make each PDA, called a unit. If you are able to sell each unit for $500, how many PDAs do you need to sell to break even?

OpenStudy (anonymous):

It just needs knowledge of basic algebra. Have you identified the variable?

OpenStudy (anonymous):

it also says Step 1: Set up the variables. Let x = the number of units y = the cost/income (expense/revenue) in dollars

OpenStudy (anonymous):

step 2 Cost Equation: total cost = fixed cost + (cost to make each unit) · (# of units) y = Income Equation: total income = y =

OpenStudy (anonymous):

step 3 Step 3: Solve the system of equations using substitution.

OpenStudy (anonymous):

step 4 Step 4: State the break-even point.

OpenStudy (anonymous):

ok so first for step 1 I set up the variables

OpenStudy (anonymous):

Let x = the number of units y = the cost/income (expense/revenue) in dollars

OpenStudy (anonymous):

x=? y=300/500(80,000/x)

OpenStudy (anonymous):

The variables have already been set up. But you dont need variable y, revenue in dollars. Just x is sufficient

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

but it says to set y up

OpenStudy (anonymous):

It can be solved more simply using just one variable.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so x=?

OpenStudy (anonymous):

number of PDAs

OpenStudy (anonymous):

which I dont know

OpenStudy (anonymous):

Which is the only thing you dont know in this equation and have to find out

OpenStudy (anonymous):

oooooo wait do I divide 80000/300?

OpenStudy (anonymous):

I dont think so. It is just the overhead cost. You have to keep it as it is.

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

so x=?

OpenStudy (anonymous):

x is number of PDAs you have to make to balance the overhead cost and the cost of making PDAs.

OpenStudy (anonymous):

ok so x=80,000+300?

OpenStudy (anonymous):

No, x will be present on both sides of the equation, which has expense amount on one side and revenue amount on the other side.

OpenStudy (anonymous):

so x=80,000+(300*x)?

OpenStudy (anonymous):

Yes that will be the left side of your equation.

OpenStudy (anonymous):

ok now what?

OpenStudy (anonymous):

Now write the right side of your equation. What will the total revenue be?

OpenStudy (anonymous):

x*y?

OpenStudy (anonymous):

y ?

OpenStudy (anonymous):

the total would be x* 80,000 * (300*x)

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

The total EXPENSE (cost of making the PDAs) will be 80000 + 300*x. I hope you understand this.

OpenStudy (anonymous):

I thought that was the left side of the equation

OpenStudy (anonymous):

Yes it was. Right side of the equation will be Money from one PDA * number of PDAs

OpenStudy (anonymous):

so x*80,00*(300*x)=500*x is my final equation?

OpenStudy (anonymous):

or no wait i would have 2 equation right? x*80,000*(300*x) and 500*x

OpenStudy (anonymous):

Not x*80,00*(300*x) , 80,000 + 300*x 80000 is not required for 1 pda, it is the total amount for all pdas

OpenStudy (anonymous):

o whoops

OpenStudy (anonymous):

so 80,000+300*x and 500*x are my equations?

OpenStudy (anonymous):

Your one equation in one variable is 80,000 + 300*x = 500*x You can solve for x easily

OpenStudy (anonymous):

An equation always has two sides

OpenStudy (anonymous):

80,000 + 300*x = 500*x 200*x=80,000 x=80,000/200 x=400

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes. Congrats! So now you are a maths whiz.

OpenStudy (anonymous):

haha thatnks but how do I find the break even point?

OpenStudy (anonymous):

The break even point is the number of PDAs you need to sell to make expense and revenue even. So it is 400, the answer you got.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

but then it says The following graph represents the above information about your hypothetical business. Label the y-axis with appropriate units and the break-even point with coordinates. Also, write on each line either "cost" or "revenue" to indicate what each represents.

OpenStudy (anonymous):

so does it want me to graph it?

OpenStudy (anonymous):

Here you have to variable y.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In the graph the x axis will be number of pdas and y axis will be expense/revenue.

OpenStudy (anonymous):

so would it look like

OpenStudy (anonymous):

wait they want coordinates? where do I put them?

OpenStudy (anonymous):